Assume fixed costs are $200/day and labor costs are $80 per day per worker (assume labor costs are the only variable costs). If the company could sell cases of snack cakes for $14 per case, how many cases of snack cakes should the company produce in order achieve the highest profit possible?

# of Workers

Q of Snack Cakes TVC TFC TC MC
0 0
1 5
1.6 10
2 15
2.35 20
2.75 25
3.22 30
3.75 35
4.35 40
5 45
5.75 50
6.6 55
7.5 60
8.5 65
10 70

Supporting Mateials

5
10
15
20
25
30
35
40
45
50
55
60
65

Question 27

Assume fixed costs are $200/day and labor costs are $80 per day per worker (assume this represents all variable costs). If the company could sell snack cakes for $14 per case, the firm would

# of Workers

Q of Snack Cakes TVC TFC TC MC
0 0
1 5
1.6 10
2 15
2.35 20
2.75 25
3.22 30
3.75 35
4.35 40
5 45
5.75 50
6.6 55
7.5 60
8.5 65
10 70

Supporting Mateials

Earn a profit
Just break even
Make a loss but produce in the short run
make a loss and produce zero output
Answer #1

Answer

TVC=number of workers *wage

TVC(Q)=w*L

w=wage and L=number of workers

TVC(5)=1*80=80

TVC(10)=1.6*80=128 and so on

TFC=$200 given

Total cost =TVC+TFC

TC(5)=80+200=280

TC(10)=128+200=328 and so on

MC(n)=(TC(n)-TC(p))/(n-p)
MC(n)=marginal cost of n th unit
TC(n)=Total cost of n units of output
TC(p)=Total cost of p unit of output
here, n>p.
MC(5)=(280-200)/(5-0)=16

MC(10)=(328-280)/(10-5)=9.6 and so on

----------

TR=P*Q

TR(5)=5*14=70

TR(10)=10*14=140 and so on

------

Profit=TR-TC

Profit(0)=0-200=-200

Profit(5)=70-280=-210 and so on

# of Workers Q of Snack Cakes TVC TFC TC MC TR Profit
0 0 0 200 200 0 -200
1 5 80 200 280 16 70 -210
1.6 10 128 200 328 9.6 140 -188
2 15 160 200 360 6.4 210 -150
2.35 20 188 200 388 5.6 280 -108
2.75 25 220 200 420 6.4 350 -70
3.22 30 257.6 200 457.6 7.52 420 -37.6
3.75 35 300 200 500 8.48 490 -10
4.35 40 348 200 548 9.6 560 12
5 45 400 200 600 10.4 630 30
5.75 50 460 200 660 12 700 40
6.6 55 528 200 728 13.6 770 42
7.5 60 600 200 800 14.4 840 40
8.5 65 680 200 880 16 910 30
10 70 800 200 1000 24 980 -20

=======================================================================

Q1

Answer
The profit is maximum at P=MC or the nearest lower MC
P=$14 and MC=$13.6 at Q=55 units
so the profit maximizing output is 55 units

-----------------
Q2
ANswer
Option 1
   Earn a profit

Profit =TR-TC
TR=770
TC=728
Profit=770-728
=42

the firm makes a profit of $42